# SweepEvGS: Event-Based 3D Gaussian Splatting for Macro and Micro Radiance Field Rendering from a Single Sweep

### v1

@@ IV-G2 How Event Noise Affects Radiance Field Reconstruction
In our experiments, we compared the performance of our method with and without the noise filter. Qualitatively, the input event is more clean on noncritical regions, indicated by Table 8. An input event stream with less noise leads to better reconstruction of radiance field representations. The quantitative results are summarized in Table IV. The metrics used for evaluation are the Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM), both of which are standard measures for assessing image quality. The results indicate that incorporating a noise filter significantly improves the rendering performance. For synthetic data, the PSNR increased from 27.8 to 29.4, and the SSIM improved from 0.90 to 0.94. Similarly, the PSNR increased from 26.9 to 29.2 for real-world data, and the SSIM improved from 0.88 to 0.91. These improvements highlight the effectiveness of our approach in enhancing the quality of the rendered radiance fields, thereby making our method more robust across different types of data.
TABLE IV: Impact of noise filtering on the rendering performance for synthetic and real-world data. Incorporating a noise filter significantly improves both PSNR and SSIM values.
@@ IV-G3 Robust Loss Function across Various Real-World Imaging Settings
Table V summarizes the quantitative comparison between our loss function and the traditional MSE loss for both synthetic and real-world data. As illustrated, our loss function outperforms the MSE loss significantly. For synthetic data, the PSNR improves from 25.2 to 29.4 and the SSIM from 0.78 to 0.94. Similarly, for real-world data, the PSNR increases from 25.7 to 29.2 and the SSIM from 0.81 to 0.91. These results highlight the effectiveness of our loss function in capturing and preserving important visual and structural details, leading to higher rendering quality compared to the traditional MSE loss. The improved performance underscores the advantages of our approach in producing more accurate and visually appealing results.
TABLE V: Quantitative comparison of rendering performance using traditional MSE loss versus our proposed loss function. Our loss function significantly improves PSNR and SSIM metrics for both synthetic and real-world data.
@@ VI Limitations and Future Work
Although SweepEvGS has shown effectiveness in novel view synthesis across different imaging setups with high efficiency, the data collection process is less adaptable. This limitation arises from the reliance on precise camera pose and extrinsic coordinate information for 3D GS methods. Currently, our real data collection can be performed on machines that can transform the event camera with a known trajectory and speed, making it possible and easy to simulate coordinates. However, an ideal industrial-level algorithm should enable sweeps to be performed by human hands, with flexibility in trajectory and speed, while still producing high-quality view synthesis and reconstruction. A straightforward approach might be to utilize event-based stereo vision to generate coordinates, thereby making the process independent of specific hardware configurations. Future work could focus on further refining this technology to achieve the goal of a flexible, high-quality view synthesis and reconstruction system that is not tied to a particular hardware setup.

### v2

@@ IV-G2 How Event Noise Affects Radiance Field Reconstruction
In our experiments, we compared the performance of our method with and without the noise filter. Qualitatively, the input event is more clean on noncritical regions, indicated by Table 8. An input event stream with less noise leads to better reconstruction of radiance field representations. The quantitative results are summarized in Table V. The metrics used for evaluation are the Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM), both of which are standard measures for assessing image quality. The results indicate that incorporating a noise filter significantly improves the rendering performance. For synthetic data, the PSNR increased from 27.8 to 29.4, and the SSIM improved from 0.90 to 0.94. Similarly, the PSNR increased from 26.9 to 29.2 for real-world data, and the SSIM improved from 0.88 to 0.91. These improvements highlight the effectiveness of our approach in enhancing the quality of the rendered radiance fields, thereby making our method more robust across different types of data.
TABLE V: Impact of noise filtering on the rendering performance for synthetic and real-world data. Incorporating a noise filter significantly improves both PSNR and SSIM values.
@@ IV-G3 Robust Loss Function across Various Real-World Imaging Settings
Table VI summarizes the quantitative comparison between our loss function and the traditional MSE loss for both synthetic and real-world data. As illustrated, our loss function outperforms the MSE loss significantly. For synthetic data, the PSNR improves from 25.2 to 29.4 and the SSIM from 0.78 to 0.94. Similarly, for real-world data, the PSNR increases from 25.7 to 29.2 and the SSIM from 0.81 to 0.91. These results highlight the effectiveness of our loss function in capturing and preserving important visual and structural details, leading to higher rendering quality compared to the traditional MSE loss. The improved performance underscores the advantages of our approach in producing more accurate and visually appealing results.
TABLE VI: Quantitative comparison of rendering performance using traditional MSE loss versus our proposed loss function. Our loss function significantly improves PSNR and SSIM metrics for both synthetic and real-world data.
@@ VI Limitations and Future Work
Although SweepEvGS has shown effectiveness in novel view synthesis across different imaging setups with high efficiency, the data collection process is less adaptable. This limitation arises from the reliance on precise camera pose and extrinsic coordinate information for 3D-GS methods. Currently, our real data collection can be performed on machines that can transform the event camera with a known trajectory and speed, making it possible and easy to simulate coordinates. However, an ideal industrial-level algorithm should enable sweeps to be performed by human hands, with flexibility in trajectory and speed, while still producing high-quality view synthesis and reconstruction. A straightforward approach might be to utilize event-based stereo vision to generate coordinates, thereby making the process independent of specific hardware configurations. Future work could focus on further refining this technology to achieve the goal of a flexible, high-quality view synthesis and reconstruction system that is not tied to a particular hardware setup.
VII Acknowledgments
This work is supported by the Theme-based Research Scheme (Grant No. T45-701/22-R), National Natural Science Foundation of China (Grant No. 62136001, 62088102), Beijing Natural Science Foundation (Grant No. L233024), and Beijing Municipal Science & Technology Commission, Administrative Commission of Zhongguancun Science Park (Grant No. Z241100003524012).